Find values for m and b in the following system so that the solution to the system is (-3,4).
5x + 7y = b
mx + y = 22

Answer by ptaylor(2198)


Our approach will be to plug the solution into the equations and solve for b and m:
5(-3)+7(4)=b
-15+28=b
b=13
and:
m(-3)+4=22 subtract 4 from both sides
-3m=22-4
-3m=18 divide both sides by -3
m=-6
So our equations are:
5x+7y=13
-6x+y=22
You can verify that (-3,4) is the solution

Hope this helps----ptaylor